8. Suppose current spot price of oil is 895 per barrel, the six-month forward price of oil is $94 per barrel, and the current risk-free rate for 6 months is 2.00% per annum. What effective six-month borrowing rate for oil does this imply? Describe the actions you would take to achieve this borrowing rate and demonstrate that these actions result in a borrowing at this rate.
Answer #1

effective 6 month interest rate r

=

n= number of compindings

r= (1+r/n)n-1=(1+0.02/2)2'-1=0.0201 or 2%

borrowings, either already made or planned, or

(ii) how much interest they might earn on deposits, either already made or planned.

If the business does not know its future interest payments or earnings, then it cannot complete a cash flow forecast accurately. It will have less confidence in its project appraisal decisions because changes in interest rates may alter the weighted average cost of capital and the outcome of net present value calculations.

There is, of course, always a risk that if a business had committed itself to variable rate borrowings when interest rates were low, a rise in interest rates might not be sustainable by the business and then liquidation becomes a possibility.

Note carefully that the primary aim of interest rate risk management (and indeed foreign currency risk management) is not to guarantee a business the best possible outcome, such as the lowest interest rate it would ever have to pay. The primary aim is to limit the uncertainty for the business so that it can plan with greater confidence.

Traditional and basic approaches

Matching and smoothing

When taking out a loan or depositing money, businesses will often have a choice of variable or fixed rates of interest. Variable rates are sometimes known as floating rates and they are usually set with reference to a benchmark such as LIBOR, the London Interbank Offered Rate. For example, variable rate might be set at LIBOR +3%.

If fixed rates are available then there is no risk from interest rate increases: a $2m loan at a fixed interest rate of 5% per year will cost $100,000 per year. Although a fixed interest loan would protect a business from interest rates increases, it will not allow the business to benefit from interest rates decreases and a business could find itself locked into high interest costs when interest rates are falling and thereby losing competitive advantage.

Similarly if a fixed rate deposit were made a business could be locked into disappointing returns.

Smoothing

In this simple approach to interest rate risk management the loans or deposits are simply divided so that some are fixed rate and some are variable rate. Looking at borrowings, if interest rates rise, only the variable rate loans will cost more and this will have less impact than if all borrowings had been at variable rate. Deposits can be similarly smoothed.

There is no particular science about this. The business would look at what it could afford, its assessment of interest rate movements and divide its loans or deposits as it thought best.

Matching

This approach requires a business to have both assets and liabilities with the same kind of interest rate. The closer the two amounts the better.

For example, let’s say that the deposit rate of interest is LIBOR + 1% and the borrowing rate is LIBOR + 4%, and that $500,000 is deposited and $520,000 borrowed. Assume that LIBOR is currently 3%.